Green Crab Experiment Part 7

Treatment Days 11 through 14

Day 11

  • Checked ammonia concentrations in all tanks
    • T1: 1.0
    • T2: 0
    • T3: 8.0
    • T4: 1.0
    • T5: 0.25
    • T6: 8.0
  • Cleaned all tanks, performed water change for all tanks, and added 1/2 tsp conditioner to T1, T3, T4, and T6
  • 1 dead crab in T3 (female green)
  • Fed ambient and warm tanks 1/4 tsp, cold tanks 1/8 tsp
  • yrv_3, yrv_9, yrv_10 still recording data (saw red flashes)
  • T3 and T6 temperatures similar (30-31ºC) when walking in and leaving on front control boxes. T6 back control box reading ~28ºC…need to borrow or order a replacement heater if temperature does not rise by tomorrow
  • Time-to-right for seven crabs in each tank

Day 12

  • Checked ammonia concentrations in all tanks
    • T1: 1.0
    • T2: 1.0
    • T3: 8.0
    • T4: 1.0
    • T5: 1.0
    • T6: 4.0
  • Cleaned all tanks, performed water change for T3 and T6, and added 1/2 tsp conditioner to all tanks
    • T6: low flow on aquarium filter, repositioned filter cartridge to improve flow
  • Fed ambient and warm tanks 1/4 tsp, cold tanks 1/8 tsp
  • Obtained HOBO data
    • yrv_3, yrv_4, and yrv_9 fine, yrv_10 low battery
    • T1 holding 12-13ºC, T4 holding 13.5ºC
    • T3 holding 31.5ºC, T6 holding 32.5ºC. Reduced T6 front control set point to 32ºC
  • T3: observed copulation/mate guarding behavior

Day 13

  • Checked ammonia concentrations in all tanks
    • T1: 0.5
    • T2: 0.25
    • T3: 8.0
    • T4: 1.0
    • T5: 0.25
    • T6: 4.0
    • Added 1/2 tsp conditioner when ammonia ≥ 1.0
  • Cleaned all tanks, performed water change for T3 and T6
  • Fed ambient and warm tanks 1/4 tsp, cold tanks 1/8 tsp
  • T6: 2 on top
  • T2: 1 dead, M YG

Day 14

  • Checked ammonia concentrations in all tanks
    • T1: 1.0
    • T2: N/a
    • T3: 8.0
    • T4: 1.0
    • T5: N/A
    • T6: 4.0
    • Added 1/2 tsp conditioner when ammonia ≥ 1.0
    • Out of solution 2, couldn’t measure ammonia in cold tanks. Added 1/2 tsp conditioner just in case
  • Cleaned all tanks, performed water change for T3 and T6, and changed all filters
  • T6: 2 on top
  • T1: 1 dead, M O
  • T3: observed copulation/mate guarding behavior, back probe out of water so I shoved it back inside. Tank seemed slightly colder than T6
  • Need to prevent salting on LED light cable, it’s slightly finnicky and not turning on in the morning unless I meddle wiht it

Going forward

  1. Continue with experiment
  2. Check HOBOs and replace low battery loggers
  3. Update protocols
  4. Organize project repository and datasheets
Written on July 18, 2022